Reaction X: ½ I2 (s) + ½ Cl2 (g) → ICl (g) ∆Hºf,RXN = 18 kJ mol-1RXN, ∆Sº298 = 78 J K-1 mol-1 Reaction Y: ½ I2 (s) + ½ Br2 (l) → IBr (g) ∆Hºf,RXN = 41 kJ mol-1RXN, ∆Sº298 = 124 J K-1 mol-1 Calculate the value of the equilibrium constant, Keq, for reaction X at 25ºC.
Answer #1

Delta G0 = DeltaH0 – T deltaS0

For reaction X; Delta G0 = 18 kJ mol-1 – (298.15K) × 78 J K-1 mol-1

For reaction X; Delta G0 = 18- 23255.7 = -23237.7 Jmol-1

Delta G0 = -RT ln Keq

Substituting Delta G0, R = 8.314 J mol-1 K-1, T= 298.15K values in above equation and solve it for Keq

-23237.7 Jmol-1 = - (8.314 J mol-1 K-1) × (298.15K) × ln Keq

ln Keq=   -23237.7 Jmol-1 / - (8.314 J mol-1 K-1) × (298.15K)

ln Keq=   9.3792

taking antilog on both sides

For reaction X ; Keq= Antilog 9.3792 = 2.344 × 10-9